How long will it take to travel one mile @ 25 MPH?
Answer by mananth(16949)

25 miles it covers in 1 hour
1 mile it travel in 1*1/25 hours
=1/25 hours or 1/25 * 60= 2.4 minute
OR
2.4 *60= 144 seconds